# PAD: Phase-Amplitude Decoupling Fusion for Multi-Modal Land Cover Classification

### v2

@@ III-B Overall Framework
As illustrated in Fig. 2, the proposed PAD framework integrates SAR and RGB modalities via a dual-branch encoder and a shared decoder, with its primary innovation in cross-modal feature fusion through phase-amplitude decoupling. The encoder employs asymmetric convolutional backbones $\mathcal{B}_{\mathrm{SAR}}$ and $\mathcal{B}_{\mathrm{RGB}}$, each with independent parameters to preserve modality-specific attributes. Specifically, $\mathcal{B}_{\mathrm{SAR}}$ extracts microwave scattering features from SAR images, while $\mathcal{B}_{\mathrm{RGB}}$ captures optical reflectance patterns from RGB images. Both backbones produce hierarchical feature maps at four stages with resolutions $\{1/4,1/8,1/16,1/32\}$ for $i\in\{0,1,2,3\}$:
$x_{RGB}^{i}=\mathcal{B}_{\mathrm{RGB}}^{(i)}\left(I_{\mathrm{RGB}}^{i}\right),\quad x_{SAR}^{i}=\mathcal{B}_{\mathrm{SAR}}^{(i)}\left(I_{\mathrm{SAR}}^{i}\right).$ (4)
Although frequency-domain fusion is more robust to modality differences, it can discard rich local interactions in the spatial domain. Therefore, we first introduce a lightweight SCF module to retain fine-grained spatial priors. The SCF module (see Eq. 5) processes the bimodal features by computing pixel-wise fusion weights via spatial attention gating. This operation preliminarily integrates complementary features in the spatial domain and reduces feature dimensionality, constraining the solution space for subsequent frequency-domain operations:
$X^{i}=\mathrm{SCF}\left(x_{RGB}^{i},x_{SAR}^{i}\right),$ (5a)
$\mathcal{S}=\mathrm{Sigmoid}\left(\mathrm{Conv}_{1\times 1}\left(\mathrm{Concat}\left(x_{1},x_{2}\right)\right)\right),$ (5b)
@@
$x=\mathrm{Conv}_{1\times 1}\left(\mathrm{Concat}\left(x_{1}^{\prime},x_{2}^{\prime}\right)\right).$ (5d)
@@
The fused feature $X^{i}$ is then transformed via FD into separate amplitude $\mathcal{A}^{i}$ and phase $\mathcal{P}^{i}$ spectra:
$\mathcal{A}^{i},\mathcal{P}^{i}=\mathrm{FD}\left(X^{i}\right).$ (6)
@@
The amplitude spectrum $\mathcal{A}^{i}$ is enhanced by the ASF module, which uses a trainable frequency radius $r$ to distinguish and reinforce high-frequency details while preserving low-frequency structures:
$\mathcal{A^{\prime}}^{i}=\mathrm{ASF}\left(\mathcal{A}^{i};r\right),$ (7)
where the PSC module applies multiplicative scaling to correct the phase:
$\mathcal{P^{\prime}}^{i}=\mathrm{PSC}\left(\mathcal{P}^{i}\right).$ (8)
Finally, FR reconstructs the fused features:
$\bm{F^{\prime}}^{i}=\mathrm{FR}\left(\mathcal{A^{\prime}}^{i},\mathcal{P^{\prime}}^{i}\right).$ (9)
@@
The shared decoder $\mathcal{D}$ then aggregates the multi-stage fused features to generate high-precision land-cover segmentation maps:
$\bm{\hat{Y}}=\mathcal{D}\left(\bm{F^{\prime}}^{0},\bm{F^{\prime}}^{1},\bm{F^{\prime}}^{2},\bm{F^{\prime}}^{3}\right).$ (10)
III-C Frequency Decoupling & Frequency Recoupling
@@
To achieve physically disentangled representations in the spectral domain, we propose a bidirectional transformation framework comprising FD and its inverse, FR. First, the input image $X\in\mathbb{R}^{H\times W}$ is transformed from the spatial domain to the frequency domain using a 2D real-input fast Fourier transform (rFFT) [48], yielding the spectrum $\mathcal{F}\in\mathbb{C}^{H\times(\lfloor W/2\rfloor+1)}$:
$\mathcal{F}\left(u,v\right)=\sum_{y=0}^{H-1}\sum_{x=0}^{W-1}X\left(y,x\right)\exp\left(-i2\pi\left(\frac{ux}{W}+\frac{vy}{H}\right)\right),$ (11)
@@
where $(u,v)$ are frequency-domain coordinates. Exploiting the conjugate symmetry of real-valued signals, we retain only the non-redundant region ($0\leq u\leq\lfloor W/2\rfloor$, $0\leq v<H$). Next, an FFT shift relocates low-frequency components to the spectrum center:
$\mathcal{F}_{\mathrm{shift}}\left(u,v\right)=\mathcal{F}\left(u\oplus\frac{W}{2},v\oplus\frac{H}{2}\right),$ (12)
where $\oplus$ denotes modulo addition. Finally, we derive the amplitude and phase spectra as
$\mathcal{A}=|\mathcal{F}|,\quad\mathcal{P}=\angle\mathcal{F},$ (13)
where $\mathcal{A},\mathcal{P}\in\mathbb{R}^{H\times(\lfloor W/2\rfloor+1)}$ capturing radiometric and structural information, respectively. This decoupling allows the model to focus on modality-specific representations: SAR phase carries scattering patterns sensitive to physical surface characteristics, whereas RGB amplitude captures fine-grained textures and semantic content. By separating them, the network avoids early-stage interference between heterogeneous modalities, enhancing feature discriminability before integration.
III-D Phase Spectrum Correction Module
@@
While previous works have assumed perfect phase alignment [17], our spectral analysis (see Fig. 1(e)) reveals that low-frequency geometric inconsistencies persist, necessitating a learnable, spatially aware correction mechanism. Inspired by channel attention, we propose a lightweight Phase Spectrum Correction (PSC) module tailored for frequency-domain phase alignment.
@@
As illustrated in Fig. 1(e), minor misalignments in the low-frequency phase spectra between SAR and RGB, attributable to geometric distortions (e.g., side-looking SAR vs. RGB orthographic projection), can compromise spatial consistency and boundary precision. Traditional rigid registration methods are inadequate for such non-rigid distortions. Thus, the PSC module performs adaptive geometric correction using learnable multiplicative masks.
@@
Unlike the global channel scaling in traditional attention mechanisms, the proposed PSC estimates a spatially varying modulation mask via a stack of $d$ consecutive $1\times 1$ convolutions, enabling localized, fine-grained phase adjustment. Given an input phase-spectrum tensor $P\in\mathbb{R}^{C\times H\times(\lfloor W/2\rfloor+1)}$, we first compress the channel dimension via $d$ $1\times 1$ convolution (reducing $C$ to $C/r$, $r\in\{4,8\}$) followed by leaky linear rectification function (Leaky ReLU) [49]. A second $1\times 1$ convolution produces a single-channel Phase Correction Mask (PCM), constrained to $(0,1)$ by a Sigmoid:
@@
$\mathrm{PCM}=\sigma\left(\mathrm{Conv}_{d}^{1\times 1}\left(\delta\left(\cdots\delta\left(\mathrm{Conv}_{1}^{1\times 1}\left(P\right)\right)\cdots\right)\right)\right),$ (14)
@@
where $d$ denotes the number of stacked convolutions and $\delta(\cdot)$ is Leaky ReLU. Pixel-wise multiplicative scaling is then applied:
$P^{\prime}=P\odot\left(1+\mathrm{PCM}\right),$ (15)
@@
where $\odot$ is the Hadamard product. This formulation enables localized, fine-grained phase adjustment while preserving phase periodicity and incurs minimal computational overhead.
III-E Amplitude Spectrum Fusion Module
@@
Previous frequency-domain methods often apply fixed high- or low-pass filters or spatially uniform spectral attention, failing to adapt to content-specific frequency distributions. In contrast, our analysis in Fig. 1(d) shows that the spectral discrepancy between SAR and RGB modalities is primarily concentrated in the medium-to-high frequency components of the amplitude spectra. To exploit this prior while maintaining gradient stability, we design the ASF module to employ a trainable frequency radius for dynamic separation and enhancement of complementary frequency bands via three sequential operations.
First, we compute a normalized frequency distance metric for each spectral component $i$:
$\mathrm{Dist}(i)=\frac{\sqrt{\left(X_{i}-X_{\mathrm{center}}\right)^{2}+\left(Y_{i}-Y_{\mathrm{center}}\right)^{2}}}{\sqrt{X_{\mathrm{max}}^{2}+Y_{\mathrm{max}}^{2}}},$ (16)
where quantifies component positions relative to the spectrum center $(X_{\mathrm{center}},Y_{\mathrm{center}})$. A trainable boundary radius $r$, constrained via a Sigmoid activation to yield $R=\sigma(r)\in[0,1]$, dynamically partitions the spectrum into low- and high-frequency regions. This constraint matches the normalized spectral dimensions and prevents $R$ from diverging to unreasonable values during training, which would otherwise lead to unstable partitioning and degraded performance.
To maintain differentiability, we employ a soft partition:
$\mathrm{Mask}_{\mathrm{high-freq}}=\sigma\left(\left(\mathrm{Dist}-R\right)\cdot\tau\right),$ (17)
@@
where temperature $\tau=10$. High-frequency components are enhanced via a position-wise MLP:
$A_{\mathrm{high-freq}}=\mathrm{MLP}\left(A\odot\mathrm{Mask}_{\mathrm{high-freq}}\right),$ (18)
@@
then concatenated with the original amplitude $A$ and fused by a $1\times 1$ convolution and residual connection:
@@
$A^{\prime}=\mathrm{Conv}_{1\times 1}\left(\mathrm{Concat}\left(A,A_{\mathrm{high-freq}}\right)\right),\quad A\leftarrow A+A^{\prime}.$ (19)
@@
This module preserves low-frequency structures while adaptively reinforcing high-frequency details that capture complementary semantic information. The learnable radius $R$ self-adjusts to input characteristics, eliminating manual threshold tuning.
@@ III-F Loss Function
Our optimization framework integrates three objectives to jointly promote semantic segmentation accuracy, multi-scale feature learning, and amplitude-phase coherence.
@@ III-F1 Primary Semantic Supervision
The primary segmentation loss is defined as the standard cross-entropy loss:
@@
where $\Omega$ denotes the set of foreground pixels (excluding background label 255), $y_{i,c}\in\{0,1\}$ indicates the ground-truth class, and $\hat{p}_{i,c}$ is the predicted probability.
@@ III-F2 Auxiliary Semantic Supervision
An auxiliary segmentation head is added at the second-to-last stage ($t=2$), also using cross-entropy loss:
$\mathcal{L}_{\mathrm{aux}}=-\sum_{i\in\Omega}\sum_{c=1}^{C}y_{i,c}\log\hat{p}^{(a)}_{i,c}.$ (21)
@@ III-F3 Phase-Amplitude Decoupling Constraint
To resolve the coupling between amplitude and phase spectra and enforce consistency between fused and original amplitudes, we introduce an amplitude consistency loss that regularizes $\mathcal{A}^{\prime}_{t}$ to remain faithful to its original decoupled state $\mathcal{A}_{t}$ across feature scales:
$\mathcal{L}_{\mathrm{amp}}=\frac{1}{|\mathcal{T}|}\sum_{t\in\mathcal{T}}\alpha_{t}\left\|\mathcal{A}^{\prime}_{t}-\mathcal{A}_{t}\right\|_{F}^{2},$ (22)
where $\mathcal{T}=\{0,1,2,3\}$ indexes the feature scales, $\alpha_{t}=0.5+0.5\frac{t}{3}$ imposes depth-progressive weighting, and $\|\cdot\|_{F}$ is the Frobenius norm.
@@ III-F4 Composite Objective
with $\lambda_{1}=0.4$ balancing auxiliary supervision and $\lambda_{2}=0.1$ controlling the amplitude regularization strength. This composite loss promotes both geometric invariance, by stabilizing phase-sensitive features through $\mathcal{L}_{\mathrm{amp}}$, and semantic invariance via hierarchical amplitude refinement.
@@ IV Experimental Results
IV-A Dataset Description
WHU-OPT-SAR [28] is a multi-modal dataset for land use classification in Hubei Province, China. It contains 100 co-registered RGB, NIR, and SAR image triples with corresponding land type labels. RGB images are acquired by GF-1, and SAR images by GF-3; both are resampled to 5 m resolution and sized 5556 $\times$ 3704 pixels. Seven primary land use categories are covered: farmland, urban areas, villages, water bodies, forests, roads, and others. We crop each image into non-overlapping 512 $\times$ 512 patches, yielding 7,000 patches, of which 80% (5,600) are used for training and 20% (1,400) for testing.
DDHR-SK is a subset of the DDHR multi-modal dataset for land cover classification in Pohang City, South Korea [29]. RGB images from GF-2 and SAR images from GF-3 are paired, and RGB images are processed with a fog filter in GIMP to simulate adverse conditions. The dataset comprises 6,173 cloud-affected RGB-SAR pairs, resampled to 1 m resolution and resized to 256 $\times$ 256 pixels. It includes five land cover types: buildings, farmland, greenery, water, and roads, with 3,087 training and 3,086 validation images in the public split.
IV-B Training Setting
All models were trained using the MMSegmentation framework [50] on a single NVIDIA H800 GPU (comparisons used an NVIDIA A100). We used a batch size of 4 for 80 k iterations, validating every 8 k iterations. Optimization was performed with AdamW [51], an initial learning rate of $1\times 10^{-4}$, and a cosine annealing schedule (T_max = total iterations). A fixed seed (42) ensured reproducibility.
We report mean IoU (mIoU), Kappa coefficient, overall accuracy (OA), and F1 score, selecting the iteration with the best mIoU. To mitigate class-imbalance bias, we compute the Kappa coefficient per sample and report its mean:
$\kappa_{i}=\frac{P_{o}^{(i)}-P_{e}^{(i)}}{1-P_{e}^{(i)}},\quad\mathrm{mKappa}=\frac{1}{M}\sum_{i=1}^{M}\kappa_{i},$ (24)
where $P_{o}^{(i)}$ and $P_{e}^{(i)}$ are the observed and expected agreements for sample $i$.
IV-C Experimental Results on Datasets
To assess the effectiveness of PAD, we replicated the original configurations of 17 comparison models, nine state-of-the-art multi-modal segmentation approaches and eight leading unimodal methods, on both WHU-OPT-SAR and DDHR-SK.
TABLE II: Quantitative Results (%) on WHU-OPT-SAR. We report per-category IoU for seven land cover classes (Farmland, City, Village, Water, Forest, Road, Others), along with OA, mKappa, mF1, and mIoU. Bold and underlined values indicate the best and second-best performance, respectively.
Method Venue IoU per Category OA mKappa mF1 mIoU

### v3

@@ III-B Overall Framework
As illustrated in Fig. 2, PAD integrates SAR and RGB via a dual-branch encoder and a shared decoder, with its primary innovation lying in cross-modal feature fusion through phase-amplitude decoupling. The encoder employs asymmetric convolutional backbones $\mathcal{B}_{\mathrm{SAR}}$ and $\mathcal{B}_{\mathrm{RGB}}$, each with independent parameters to preserve modality-complementary attributes. Specifically, $\mathcal{B}_{\mathrm{SAR}}$ extracts microwave backscattering features from SAR images, while $\mathcal{B}_{\mathrm{RGB}}$ captures optical reflectance patterns from RGB images. Both backbones produce hierarchical features at four stages with resolutions $\{1/4,1/8,1/16,1/32\}$ for $i\in\{0,1,2,3\}$:
$x_{\mathrm{RGB}}^{i}=\mathcal{B}_{\mathrm{RGB}}^{(i)}\left(I_{\mathrm{RGB}}\right),\quad x_{\mathrm{SAR}}^{i}=\mathcal{B}_{\mathrm{SAR}}^{(i)}\left(I_{\mathrm{SAR}}\right).$ (4)
Although frequency-domain fusion is robust to modality differences, it may discard rich local interactions. To retain fine-grained spatial priors, we introduce a lightweight SCF module. SCF computes pixel-wise fusion weights via spatial attention gating and reduces feature dimensionality to constrain the solution space for subsequent spectral operations:
$X^{i}$ $=\mathrm{SCF}\left(x_{\mathrm{RGB}}^{i},x_{\mathrm{SAR}}^{i}\right),$ (5a)
$\mathcal{S}$ $=\sigma\left(\mathrm{Conv}_{1\times 1}\left(\mathrm{Concat}\left(x_{1},x_{2}\right)\right)\right),$ (5b)
@@
$x$ $=\mathrm{Conv}_{1\times 1}\left(\mathrm{Concat}\left(x_{1}^{\prime},x_{2}^{\prime}\right)\right),$ (5d)
where $(x_{1},x_{2})=(x_{\mathrm{RGB}}^{i},x_{\mathrm{SAR}}^{i})$; $\sigma(\cdot)$ is the Sigmoid; $\mathrm{Conv}_{1\times 1}$ denotes a $1\times 1$ convolution; $\mathrm{Concat}(\cdot,\cdot)$ concatenates along channels; and $\odot$ denotes the Hadamard (element-wise) product.
@@
The fused feature $X^{i}$ is then transformed to separate amplitude $\mathcal{A}^{i}$ and phase $\mathcal{P}^{i}$ spectra:
$\mathcal{A}^{i},\mathcal{P}^{i}=\mathrm{FD}\left(X^{i}\right),$ (6)
@@
where $\mathrm{FD}(\cdot)$ denotes the frequency decoupling operator (defined in §III-C). ASF enhances the amplitude spectrum using a trainable frequency radius $r$ to distinguish and reinforce high-frequency details while preserving low-frequency structures:
$\mathcal{A}^{\prime i}=\mathrm{ASF}\left(\mathcal{A}^{i};r\right).$ (7)
PSC applies multiplicative scaling to correct the phase:
$\mathcal{P}^{\prime i}=\mathrm{PSC}\left(\mathcal{P}^{i}\right).$ (8)
Finally, inverse reconstruction produces fused features:
$\bm{F}^{\prime i}=\mathrm{FR}\left(\mathcal{A}^{\prime i},\mathcal{P}^{\prime i}\right),$ (9)
@@
where $\mathrm{FR}(\cdot,\cdot)$ is the inverse frequency recoupling operator (defined in §III-C). A shared decoder $\mathcal{D}$ aggregates multi-stage fused features to generate the land-cover map:
$\widehat{\bm{Y}}=\mathcal{D}\left(\bm{F}^{\prime 0},\bm{F}^{\prime 1},\bm{F}^{\prime 2},\bm{F}^{\prime 3}\right).$ (10)
III-C Frequency Decoupling and Recoupling
@@
To obtain physically decoupled spectral representations, we employ a bidirectional transformation comprising frequency decoupling (FD) and inverse recoupling (FR). Given $X\in\mathbb{R}^{H\times W}$, a 2D real-input fast Fourier transform (rFFT) [52] yields:
$\mathcal{F}(u,v)=\sum_{y=0}^{H-1}\sum_{x=0}^{W-1}X(y,x)\exp\left(-\mathrm{i}2\pi\left(\frac{ux}{W}+\frac{vy}{H}\right)\right),$ (11)
@@
where $(u,v)$ are frequency coordinates. Exploiting conjugate symmetry for real inputs, we retain only the non-redundant region and center low-frequency components via an FFT shift:
$\mathcal{F}_{\mathrm{shift}}(u,v)=\mathcal{F}\left(u\oplus\frac{W}{2},v\oplus\frac{H}{2}\right),$ (12)
where $\oplus$ is modulo addition. Amplitude and phase are:
$\mathcal{A}=\left|\mathcal{F}_{\mathrm{shift}}\right|,\quad\mathcal{P}=\angle\mathcal{F}_{\mathrm{shift}},$ (13)
with $\mathcal{A},\mathcal{P}\in\mathbb{R}^{H\times(\lfloor W/2\rfloor+1)}$. FR applies the inverse operations to reconstruct spatial features from $(\mathcal{A},\mathcal{P})$.
III-D Phase Spectrum Correction (PSC) Module
@@
While prior works often assume perfect phase alignment [18], our spectral analysis (see Fig. 1(e)) shows that low-frequency geometric inconsistencies persist, necessitating a learnable, spatially aware correction mechanism. Inspired by channel attention, we propose a lightweight Phase Spectrum Correction (PSC) module tailored for frequency-domain phase alignment.
@@
As shown in Fig. 1(e), minor misalignments in the low-frequency phase spectra between SAR and RGB, attributable to geometric distortions (e.g., side-looking SAR vs. orthographic RGB), can compromise spatial consistency and boundary precision. Traditional rigid registration is inadequate for such non-rigid distortions. PSC performs adaptive geometric correction using learnable multiplicative masks.
@@
Unlike global channel scaling in traditional attention, PSC estimates a spatially varying modulation mask via a stack of $d$ consecutive $1{\times}1$ convolutions, enabling localized, fine-grained phase adjustment. Given an input phase-spectrum tensor $P\in\mathbb{R}^{C\times H\times(\lfloor W/2\rfloor+1)}$, we first compress channels with $d$ successive $1{\times}1$ convolutions (reducing $C$ to $C/r$, $r\in\{4,8\}$), each followed by a Leaky ReLU [53]. A final $1{\times}1$ convolution produces a single-channel Phase Correction Mask (PCM), constrained to $(0,1)$ by a Sigmoid:
@@
$\mathrm{PCM}=\sigma\left(\mathrm{Conv}^{1\times 1}_{d}\left(\delta\left(\cdots\delta\left(\mathrm{Conv}^{1\times 1}_{1}\left(P\right)\right)\cdots\right)\right)\right),$ (14)
@@
where $\delta(\cdot)$ is Leaky ReLU and $\mathrm{Conv}^{1\times 1}_{k}$ denotes the $k$-th $1{\times}1$ convolution in the stack. Pixel-wise multiplicative scaling is then applied:
$P^{\prime}=P\odot\left(1+\mathrm{PCM}\right).$ (15)
@@
This formulation enables localized, fine-grained phase adjustment while preserving phase periodicity and incurring minimal computational overhead.
III-E Amplitude Spectrum Fusion (ASF) Module
@@
Previous frequency-domain methods often employ fixed high-/low-pass filters or spatially uniform spectral attention, which cannot adapt to content-specific frequency distributions. In contrast, our analysis in Fig. 1(d) shows that RGB-SAR spectral discrepancies are concentrated in medium-to-high amplitude frequencies. To exploit this prior while maintaining gradient stability, ASF uses a trainable frequency radius to dynamically separate and enhance complementary bands via three steps. First, for each spectral component indexed by $i$, we compute a normalized radial distance:
$\mathrm{Dist}(i)=\frac{\sqrt{\left(X_{i}-X_{\mathrm{c}}\right)^{2}+\left(Y_{i}-Y_{\mathrm{c}}\right)^{2}}}{\sqrt{X_{\max}^{2}+Y_{\max}^{2}}},$ (16)
where $(X_{\mathrm{c}},Y_{\mathrm{c}})$ is the spectrum center and $(X_{\max},Y_{\max})$ are the maximal coordinate indices along each axis, so $\mathrm{Dist}(i)\in[0,1]$ quantifies each component’s position relative to the center. A trainable boundary radius $r$ is mapped by a Sigmoid to $R=\sigma(r)\in[0,1]$, partitioning the spectrum into low-/high-frequency regions and preventing divergence during training. To maintain differentiability, we use a soft mask:
$\mathrm{Mask}_{\mathrm{hi}}=\sigma\left(\left(\mathrm{Dist}-R\right)\cdot\tau\right),$ (17)
@@
where $\tau=10$ is a temperature hyperparameter. High-frequency components are enhanced via a position-wise MLP:
$A_{\mathrm{hi}}=\mathrm{MLP}\left(A\odot\mathrm{Mask}_{\mathrm{hi}}\right),$ (18)
@@
and then concatenated with the original amplitude $A$ and fused by a $1{\times}1$ convolution with a residual update:
@@
$A^{\prime}=\mathrm{Conv}_{1\times 1}\left(\mathrm{Concat}\left(A,A_{\mathrm{hi}}\right)\right),\quad A\leftarrow A+A^{\prime},$ (19)
@@
where $A\in\mathbb{R}^{C\times H\times(\lfloor W/2\rfloor+1)}$ is the input amplitude spectrum, and $\mathrm{MLP}(\cdot)$ is a channel-wise two-layer perceptron with nonlinearity. This preserves low-frequency structure while adaptively reinforcing high-frequency details carrying modality-complementary semantics. The learnable radius $R$ self-adjusts to input characteristics, eliminating manual threshold tuning.
@@ III-F Loss Function
We integrate three objectives to jointly promote semantic accuracy, multi-scale representation learning, and amplitude-phase coherence.
@@ III-F1 Primary Semantic Supervision
where $\Omega$ is the set of labeled pixels (excluding the ignore label 255), $y_{i,c}\in\{0,1\}$ is the one-hot ground truth for class $c$, and $\hat{p}_{i,c}$ is the predicted probability.
@@ III-F2 Auxiliary Semantic Supervision
An auxiliary head at the penultimate stage ($t=2$) uses the same objective:
$\mathcal{L}_{\mathrm{aux}}=-\sum_{i\in\Omega}\sum_{c=1}^{C}y_{i,c}\log\hat{p}^{(a)}_{i,c}$, (21)
where $\hat{p}^{(a)}_{i,c}$ denotes auxiliary predictions.
@@ III-F3 Phase-Amplitude Decoupling Constraint
To regularize the fused amplitude towards its decoupled counterpart across scales, we define:
$\mathcal{L}_{\mathrm{amp}}=\frac{1}{\left|\mathcal{T}\right|}\sum_{t\in\mathcal{T}}\alpha_{t}\left\|\mathcal{A}^{\prime}_{t}-\mathcal{A}_{t}\right\|_{F}^{2},$ (22)
where $\mathcal{T}=\{0,1,2,3\}$ indexes feature scales; $\alpha_{t}=0.5+0.5t/3$ applies depth-progressive weighting; $\|\cdot\|_{F}$ is the Frobenius norm; and $\mathcal{A}_{t},\mathcal{A}^{\prime}_{t}$ are the original and fused amplitude spectra at scale $t$.
@@ III-F4 Composite Objective
where $\lambda_{1}=0.4$ balances auxiliary supervision and $\lambda_{2}=0.1$ controls amplitude regularization. This composite loss promotes geometric invariance by stabilizing phase-sensitive features via $\mathcal{L}_{\mathrm{amp}}$, and semantic robustness via hierarchical amplitude refinement.
@@ IV Experimental Results
IV-A Experimental Setup
IV-A1 Dataset Description
WHU-OPT-SAR [29] is a multi-modal land-use dataset covering Hubei Province, China. It contains 100 co-registered RGB-SAR-NIR triples labeled with seven categories: farmland, urban, village, water, forest, roads, and others. RGB (GF-1) and SAR (GF-3) images are resampled to 5 m resolution and sized $5556\times 3704$ pixels. Each image is cropped into non-overlapping $512\times 512$ patches, yielding 7,000 samples split 80%/20% for training/testing.
DDHR-SK [30] is a fog-degraded subset of DDHR from Pohang, South Korea, pairing GF-2 RGB and GF-3 SAR images. It contains 6,173 cloud-affected RGB-SAR pairs, resampled to 1 m and resized to $256\times 256$ pixels, with five categories: buildings, farmland, greenery, water, and roads. The public split provides 3,087 training and 3,086 validation samples.
IV-A2 Training Settings
All models are trained using the MMSegmentation framework [54] on a single NVIDIA H800 GPU (comparisons use an NVIDIA A800). We use a batch size of 4 for 80k iterations, validating every 8k iterations. Optimization is performed with AdamW, an initial learning rate of $1\times 10^{-4}$, and a cosine-annealing schedule (T_max = total iterations). A fixed seed (42) ensures reproducibility.
IV-A3 Evaluation Metrics
We report mean Intersection-over-Union (mIoU), overall accuracy (OA), F1 score, and Cohen’s Kappa (mKappa), selecting the checkpoint with the highest mIoU. To mitigate class-imbalance bias, Kappa is computed for each test sample and then averaged across all samples to obtain mKappa.
IV-B Main Quantitative Results
To assess PAD’s effectiveness, we replicate the original configurations of 20 comparison models, 12 state-of-the-art multi-modal segmentation approaches and 8 leading unimodal methods, on both WHU-OPT-SAR and DDHR-SK.
TABLE II: Quantitative Results (%) on WHU-OPT-SAR. Per-class IoU is reported for seven land-cover categories (Farmland, City, Village, Water, Forest, Road, and Others), along with OA, mKappa, mF1, and mIoU. The best and second-best results are highlighted in bold and underlined, respectively.
Method Venue Per-class IoU OA mKappa mF1 mIoU